# Are EU low-carbon structural funds efficient in reducing emissions?

### v1

@@ 3 Data and methods
We focus on emission data reported by the Emissions Database for Global Atmospheric Research (EDGAR), managed by the Joint Research Centre of the European Commission. EDGAR provides quantitative estimates of global GHG emissions by country and sector, including CO2, CH4, N2O, and F-gases, using a methodology aligned with Intergovernmental Panel on Climate Change (IPCC) guidelines (Crippa et al., 2023). Specifically, we use European regional data at the NUTS 2 level of aggregation, focusing on aggregated GHG emissions and their breakdown into CO2, CH4, and N2O emissions.
Our primary objective is to analyse the impacts on GHG emissions resulting from expenditure on the thematic objective related to the low-carbon economy in the 2007-2013 and 2014-2020 ESI funds programmes (a more detailed definition of the expenditure variables is presented below). In the econometric modelling setup, we use other relevant covariates to control for economic activity, sectoral composition, and local policy stringency that influence emissions across European regions.
@@ 3.1 Data
The total expenditures reflect the targeted efforts to reduce economic disparities within the EU. Certain regions consistently receive higher funds according to ongoing economic development needs or projects requiring sustained investment. The disparities in the distribution of funds consistently with eligibility show that less developed regions in Southern and Eastern Europe generally receive higher per capita funding than Western and Northern Europe regions.
@@
For the 2014-2020 period, the low-carbon expenditures are straightforward to determine since there is a categorical variable related to thematic objective 4 (TO-4), labelled as “low-carbon economy”. The aim of the expenditure can be tracked by the reported “intervention field” in the data. Specifically, related to TO-4, the intervention fields include renewable energy (wind/solar/biomass), energy efficiency renovation of public infrastructure, cycle tracks and footpaths, among others (see Table A1 in the Appendix).
For the 2007-2013 program, there is no categorical variable explicitly associated with the low-carbon economy as in the 2014-2020 program. However, the raw categorisation data describes the “priority” related to the reported expenditures, serving as a proxy for the aim of the expenditure. We use this information to create a low-carbon expenditures variable for this period. Priority labels related to this period include cycle tracks, intelligent transport systems, renewable energy (wind/solar/biomass/hydroelectric), energy efficiency, co-generation, energy management, and promotion of clean urban transport. The priority codes are reported in Table A1 of the Appendix.
A critical aspect of the ESI funds expenditure data is the absence of detailed annual expenditure information at the regional level. While reported expenditures can be easily tracked at the regional level, the reports correspond to the entire programming period. To our knowledge, there is no disaggregated, time-varying consolidated official information. Nevertheless, the cohesion data platform presents estimations of the expenditure by NUTS 2, which are aggregated for the most significant types of structural funds only. Therefore, to address our need for yearly data, we have opted for two approaches: one assumes that expenditures are distributed evenly over time (dividing the reported expenditures over the programming period), and the other assumes they follow the same pattern as the total aggregated expenditure reported by the cohesion data platform.444Data available at: https://cohesiondata.ec.europa.eu/Other/Historic-EU-payments-regionalised-and-modelled/tc55-7ysv/about_data. Since both strategies yielded qualitatively similar conclusions, we present in the main text the results obtained with the first approach, while those derived from the second are included in the Appendix as a robustness check (Tables A2 and A3).
@@
Economic variables were obtained from the ARDECO (the Annual Regional Database of the European Commission’s Directorate General for Regional and Urban Policy), which contains annual data at the regional level and is maintained and updated by the Joint Research Centre of the European Commission. We use variables such as the real gross domestic product (GDP), the population, and the real gross value added (GVA) for all industry sectors according to NACE Rev. 2 definitions. These sectors are A: Agriculture, Forestry and Fishing; B–E: Industry; F: Construction; G–J: Wholesale, Retail, Transport, Accommodation & Food Services, Information and Communication; K–N: Financial & Business Services; and O–U: Non-market Services. The choice of these variables is motivated by their relevance to understanding how economic activity is related to emissions. In particular, we include industry-level GVA to control direct sources of emissions and to better capture regional emission profiles, given that the participation of industries varies significantly across European regions.
@@
To assess the effects of environmental policies on emissions, we use the EPS index (OECD, 2016). This index, developed by Brunel and Levinson (2013) and further refined by Botta and Koźluk (2014), provides a standardised composite measure incorporating both market and non-market-based policies, covering 27 OECD countries over the period from 1990 to 2020.
@@
Finally, to determine the political ideology at the regional level, we calculate the regional weighted average of political ideologies using the results of European subnational election data. To measure political ideology, we use the “left/right” dimension variable reported by ParlGov (Döring et al., 2022) and complemented by CHES-Europe (Jolly et al., 2022), which takes values in the interval 0–10, with values close to zero representing extreme left ideology and values close to 10 representing extreme right ideology. We use EU-NED (Schraff et al., 2023), which reports sub-national election data across European countries over the last three decades at the party level. We assign each party at the regional level an ideology by combining information from ParlGov and CHES-Europe. These databases are integrated within the Party Facts platform, which facilitates the incorporation of party-level data.555The data is available at: https://www.chesdata.eu/ches-europe and https://www.parlgov.org/. Also see: https://partyfacts.herokuapp.com/. We include this variable to provide additional controls for political ideology. While our focus is not on analysing how politics affects emissions, this is a growing area of interest in political science. For instance, Jahn (2021) highlights regional variations in the impact of populist governance on environmental policies, showing that right-wing populist administrations in North Western and Eastern Europe are linked to heightened GHG emissions, whereas their left-wing counterparts in Southern Europe are associated with decreases in emissions.
@@
Figure 2 presents a set of maps illustrating different facets of the environmental impact and distribution of the Cohesion Funds. In the upper part, three maps show average GHG emissions in shades of blue to red, where blue denotes lower levels and red denotes higher levels. The first map, on the left, represents the average emissions for 2015-2019 in teragrams of CO2 equivalent per year (Tg CO2eq/year). The other two maps compare the average annual per capita emissions for 2000-2004 and 2015-2019, respectively. These maps show that, on average, towards the end of the 2010s, most European regions emitted less than 50 teragrams of CO2. However, there is significant heterogeneity in per capita emissions, highlighting areas of intense economic activity or low population density. Comparing the early 2000s and late 2010s, it is evident that many regions with initially high emissions tended to reduce them (see, for instance, the Castile region in Spain), while some regions maintained or increased their per capita emission levels (for instance, the Masovian region in Poland).
@@
This pattern raises questions about the effectiveness of the allocation of funds relative to emission levels. In Figure 3, we plot the total GHG emissions per capita versus the total expenditure in low-carbon economy funds per capita. Emissions are considered at the beginning of the corresponding programming period. The plot exhibits great moderation in emission levels per capita for most European regions and shows that higher funds per capita were consistently allocated to less developed regions. However, a positive correlation pattern in the allocation is less clearly seen in both panels, suggesting that these funds are not necessarily allocated to the regions with the highest need to reduce emissions.
@@
where is an emission indicator for region at time , and and are the cumulative yearly programme expenditures per capita (in log) for the periods 2007-2013 and 2014-2020, respectively, with denoting lags to acknowledge potential delayed effects. We include fixed effects for location () and time () to control for time-invariant unobserved heterogeneity across regions as well as uniform temporal shocks, such as those associated with environmental directives instituted at the European Union level. Finally, the term is the estimation residual.
In the left-hand in Eq. (1), the term serves as a reference notation for different emission variables. Specifically, we use four different dependent variables based on the total emissions of region at time . These are the levels of emissions per capita (in log), defined as GHGEpci,t=; the emission intensity relative to the GDP, defined as GHGEii,t=; and the growth rates of GHGEpci,t and GHGEii,t.
@@
This preliminary exercise demonstrates that the relationship between low-carbon expenditures and GHG levels, as well as the growth rates of emissions per capita (GHGEpc) and emission intensity (GHGEi), can be mixed. These mixed results suggest that the econometric specification might be masking underlying factors that could restrict the effectiveness of low-carbon funds. One relevant limiting factor is the persistent disparities among European regions. Indeed, the development level of regions is a key factor influencing the effectiveness of cohesion funds (Boldrin and Canova, 2001; Rodríguez-Pose and Fratesi, 2004). However, many aspects related to the heterogeneity of European regions can explain the mixed results. For instance, less developed regions might use these funds to build sustainable infrastructure to reduce emissions, while more developed regions, having already made progress in emissions efficiency, might experience a more moderate impact from these funds. Therefore, the empirical strategy must consider that short-term outcomes may differ from longer-term impacts. This consideration is crucial for environmental policy because it suggests that sustainability investments may be misallocated, leading to unintended or unfavourable effects on emissions.
@@ 3.2 Methodology
Our methodology shifts from this type of analysis, which primarily assesses the impact of GDP growth on emissions, to a comparative panel data analysis focusing on the effects of expenditure in the low-carbon economy. We specifically highlight the role of cohesion fund investments in fostering low-carbon economies to curb emissions. Therefore, it is central in our methodology to decompose time series variables, such as emissions, GDP, and GVA, into their cyclical and trend components for each region independently. Then we consider these time series in two different panel data model frameworks: regional trends and regional cycles of emissions. In this way, we analyse the difference between long-term trends and cyclical effects of the impact of low-carbon economy funds on GHG emissions.
@@
After decomposing the GHGEpc and GHGEi time series, we independently analyse the panels of regional emission trends and cycles, considering the regional development level differences. The model for emissions trends is defined as follows:
where the superscript specifies that we conduct the estimation on time series trend components. This way, is the trend of the logarithm of emissions per capita for NUTS 2 region at time , similarly is the trend component of the GDP (in log) of region at time , and and are the yearly cumulative programme expenditures (in log) at time . We include the development level of region using the dummy variable , where . The term consider additional control variables such as the dummy variables capturing the effect of stronger stringency of environmental policies, denoted by 666Following Naqvi (2021), we measure the impact of stronger environmental policies employing a dummy variable that takes the value of one when EPS index is greater or equal than 2.5., and the term , which indicates the average political ideology of the region, determined by the outcomes of local elections (see data section for details). This model incorporates fixed effects for both location () and time (), removing biases related to unobserved fixed regional heterogeneity and factors that vary over time but impact all regions equally.
Additionally to Eq. (3) focusing on the trend of regional emissions, we also consider the impact of low-carbon investments on the cyclical components of emissions using the following specification:
Note that the superscript now represents the emissions’ cyclical components, therefore, is the cycle component of the GDP (in log) of region at time .
In our specifications, Eq. (3) and (4), the parameters of interest are , , and , which capture how a region’s development level can influence structural funds’ effect in emissions components. The aim is to explore whether the impact of these expenditures varies according to the economic development level of the region, affecting the cyclical and trend components of emissions.
As a robust check, in our specifications, we replace GDP with the set of GVA variables for all economic sectors. This allows us to better control how regional productive structures affect GHG emissions.
@@ 4.1 Effect on emissions trend
Table 2 presents the estimation results of Eq (3) using as the dependent variable the trends for GHGEpc or GHGEi. The results related to the expenditures for the different groups of regions are presented with respect to the most developed regions (DER). The table reveals mixed patterns that underline the importance of regional development level in the influence of expenditure.
@@
| GHGEpc | GHGEi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.158∗∗ | 0.138∗∗ | 0.012 | 0.159∗∗ | 0.171∗∗∗ | 0.108∗∗ |
@@
| Political Ideology | 0.016∗∗∗ | 0.001 | 0.002 | 0.007∗∗ | ||

### v2

@@ 3 Data and methods
We aim to analyse the impact of expenditure on GHG emissions resulting from expenditure on the thematic objective related to the low-carbon economy in the 2007-2013 and 2014-2020 ESI funds programmes (a more detailed definition of the expenditure variables is presented below). We shall use relevant covariates to control economic activity, sectoral composition, and environmental policy stringency.
@@ 3.1 Data
Emissions data
We focus on emissions data reported by the Emissions Database for Global Atmospheric Research (EDGAR), managed by the Joint Research Centre of the European Commission. EDGAR provides quantitative estimates of global GHG emissions by country and sector, including CO2, CH4, N2O, and F-gases, using a methodology aligned with Intergovernmental Panel on Climate Change (IPCC) guidelines (Crippa et al., 2023). Specifically, we use European regional data at the NUTS 2 level of aggregation, focusing on aggregated GHG emissions and their breakdown into CO2, CH4, and N2O emissions.
@@
The total expenditures reflect the targeted efforts to reduce economic disparities within the EU. Certain regions consistently receive higher funds according to ongoing economic development needs or projects requiring sustained investment. The disparities in the distribution of funds, consistent with eligibility, show that less developed regions in Southern and Eastern Europe generally receive higher per capita funding than Western and Northern Europe regions.
@@
For the 2014-2020 programme, the low-carbon expenditures are straightforward to determine since there is a categorical variable related to thematic objective 4 (TO-4), labelled as “low-carbon economy”. The aim of the expenditure can be tracked by the reported “intervention field” in the data. Specifically, related to TO-4, the intervention fields include renewable energy (wind/solar/biomass), energy efficiency renovation of public infrastructure, cycle tracks and footpaths, among others (see Table A1 in the Appendix).
For the 2007-2013 programme, there is no categorical variable explicitly associated with the low-carbon economy as in the 2014-2020 programme. However, the raw categorisation data describe the “priority” related to the reported expenditures, serving as a proxy for the aim of the expenditure. We use this information to create a low-carbon expenditures variable for this period. Priority labels related to this period include cycle tracks, intelligent transport systems, renewable energy (wind/solar/biomass/hydroelectric), energy efficiency, co-generation, energy management, and promotion of clean urban transport. The priority codes are reported in Table A1 of the Appendix.
A critical aspect of the ESI funds expenditure data is the absence of detailed annual expenditure information at the regional level. While reported expenditures can be easily tracked at the regional level, the reports correspond to the entire programming period. To our knowledge, there is no disaggregated time-varying consolidated official information available. Given this constraint, we built an annually growing cumulative investment series for each region and programme. This implies that, within each programming period, the total per capita expenditure is allocated progressively across the corresponding years. While this introduces a smoothing assumption, it allows us to preserve the annual variation in emissions and covariates. Therefore, considering the capital-intensive nature of the funded projects, we argue that their environmental effects are plausibly associated with the cumulative stock of investments. We further validate this modelling choice through robustness checks, including alternative cross-sectional estimations presented in Appendix Tables A2 and A3.
@@
Economic variables were obtained from the ARDECO (the Annual Regional Database of the European Commission’s Directorate General for Regional and Urban Policy), which contains annual data at the regional level and is maintained and updated by the Joint Research Centre of the European Commission. We use variables such as the real gross domestic product, the population, and the real gross value added (GVA) for all industry sectors according to NACE Rev. 2 definitions. These sectors are A: Agriculture, Forestry and Fishing; B–E: Industry; F: Construction; G–J: Wholesale, Retail, Transport, Accommodation & Food Services, Information and Communication; K–N: Financial & Business Services; and O–U: Non-market Services. The choice of these variables is motivated by their relevance to understanding how economic activity is related to emissions. In particular, we include industry-level GVA to control direct sources of emissions and to better capture regional emission profiles, given that the participation of industries varies significantly across European regions.
@@
To assess the effects of environmental policies on emissions, we use the EPS index (OECD, 2016). This index, developed by Brunel and Levinson (2013) and further refined by Botta and Koźluk (2014), provides a standardised composite measure incorporating both market and non-market-based policies, covering 27 OECD countries between 1990 and 2020.
@@
Finally, to determine the political ideology at the regional level, we calculate the regional weighted average of political ideologies using the results of European subnational election data. To measure political ideology, we use the “left/right” dimension variable reported by ParlGov (Döring et al., 2022) and complemented by CHES-Europe (Jolly et al., 2022), which takes values in the interval 0-10, with values close to zero representing extreme left ideology and values close to 10 representing extreme right ideology. We use EU-NED (Schraff et al., 2023), which reports sub-national election data across European countries over the last three decades at the party level. We assign each party at the regional level an ideology by combining information from ParlGov and CHES-Europe. These databases are integrated within the Party Facts platform, which facilitates the incorporation of party-level data.444The data are available at: https://www.chesdata.eu/ches-europe and https://www.parlgov.org/. Also see: https://partyfacts.herokuapp.com/. We include this variable to provide additional controls for political ideology. While our focus is not on analysing how politics affects emissions, this is a growing area of interest in political science. For instance, Jahn (2021) highlights regional variations in the impact of populist governance on environmental policies, showing that right-wing populist administrations in North Western and Eastern Europe are linked to heightened GHG emissions, whereas their left-wing counterparts in Southern Europe are associated with decreases in emissions.
@@
Figure 2 presents a set of maps illustrating different facets of the environmental impact and distribution of the ESI funds. In the upper part, three maps show average GHG emissions in shades of blue to red, where blue denotes lower levels and red denotes higher levels. The first map, on the left, represents the average emissions for 2015-2019 in teragrams of CO2 equivalent per year (Tg CO2eq/year). The other two maps compare the average annual per capita emissions for 2000-2004 and 2015-2019, respectively. These maps show that, on average, towards the end of the 2010s, most European regions emitted less than 50 teragrams of CO2. However, there is significant heterogeneity in per capita emissions, highlighting areas of intense economic activity or low population density.555Sparsely populated regions may have high per capita emissions despite relatively moderate total emissions. This is due to emission-intensive activities, like agriculture and energy production, concentrated in some low-density areas. Comparing the early 2000s and late 2010s, it is evident that many regions with initially high emissions tended to reduce them (for instance, the Castile region in Spain), while some regions maintained or increased their per capita emission levels (for instance, the Masovian region in Poland).
@@
where is an emission indicator for region at time , and and are the cumulative yearly programme expenditures per capita for the periods 2007-2013 and 2014-2020, respectively, with denoting lags to acknowledge potential delayed effects. We include fixed effects for location () and time () to control for time-invariant unobserved heterogeneity across regions as well as uniform temporal shocks, such as those associated with environmental directives instituted at the European Union level. Finally, the term is the estimation residual.
In the left-hand side of Eq. (1), the term serves as a reference notation for different emission variables. Specifically, we use four different dependent variables based on the total emissions of region at time . These are the levels of emissions per capita (in log), defined as GHGEpci,t=; the emission intensity relative to the GDP, defined as GHGEii,t=; and the growth rates of GHGEpci,t and GHGEii,t.
@@
These mixed results suggest that the econometric specification might be masking underlying factors that could restrict the effectiveness of low-carbon funds. One relevant limiting factor is the persistent disparities among European regions. Indeed, the development level of regions is a key factor influencing the effectiveness of the ESI funds (Boldrin and Canova, 2001; Rodríguez-Pose and Fratesi, 2004). However, many aspects related to the heterogeneity of European regions can explain the mixed results. For instance, less developed regions might use these funds to build sustainable infrastructure to reduce emissions, while more developed regions, having already made progress in emissions efficiency, might experience a more moderate impact from these funds. Therefore, the empirical strategy must consider that short-term outcomes may differ from longer-term impacts. This consideration is crucial for environmental policy because it suggests that sustainability investments may be misallocated, leading to unintended or unfavourable effects on emissions.
@@ 3.2 Methodology
Our methodology shifts from this type of analysis, which primarily assesses the impact of GDP growth on emissions, to a comparative panel data analysis focusing on the effects of expenditure in the low-carbon economy. We specifically highlight the role of the ESI funds in fostering low-carbon economies to curb emissions. Therefore, it is central in our methodology to decompose time series variables, such as emissions, GDP, and GVA, into their cyclical and trend components for each region independently. Then, we consider these time series in two different panel data model frameworks: regional trends and regional cycles of emissions. In this way, we analyse the difference between long-term trends and cyclical effects of the impact of low-carbon economy funds on GHG emissions.
@@
After decomposing the GHGEpc and GHGEi time series, we independently analyse the panel data of regional emission trends and cycles, taking into account differences in regional development levels. The model for the trend and cycle components of emissions is defined as follows:
where the superscript indicates that we estimate the model separately for the trend and cycle components of the series. Specifically, denotes either the trend or the cycle component of the logarithm of an emissions indicator (GHGEpc or GHGEi)666 Using these relative measures offers several advantages: first, per‑capita emissions normalise for population size, allowing us to compare each region’s carbon footprint regardless of its scale; second, emissions intensity controls for economic scale, highlighting how efficiently regions produce output per unit of carbon. This enhances comparability across diverse regions and reduces heteroskedasticity in panel regressions. More importantly, it allows us to better isolate the impact of low‑carbon investments on environmental efficiency and avoid confounding with demographic or economic differences. for region at time , and is the corresponding component of the logarithm of GDP. The terms and refer to the cumulative yearly expenditures of the 2007–2013 and 2014–2020 programmes, respectively, lagged by years. Regional development levels are introduced via the dummy variable , with representing less developed, transition, and developed regions, respectively.
The vector includes additional controls, such as a dummy for environmental policy stringency ()777Following Naqvi (2021), we use a dummy equal to one when the EPS index is greater than or equal to 2.5. and an indicator of average political ideology (), based on local election outcomes. The model also incorporates region fixed effects () and year fixed effects () to control for time-invariant regional characteristics and common shocks across all regions. As a robustness check, we replace GDP with the set of GVA variables for all economic sectors to better account for the role of regional productive structures in explaining GHG emissions.
While our specification includes key relevant variables, we acknowledge that endogeneity may still pose concerns. For instance, one might argue that environmental regulations or improvements in energy efficiency could, in principle, influence regional output and investment decisions. For these reasons, we have included a five-year lag for all explanatory variables, ensuring that the predictors precede the observed emissions outcomes by a meaningful temporal distance. Therefore, while our specification mitigates (but does not fully eliminate) simultaneity concerns, it relies on the timing assumption that lagged expenditures impact emissions.
Moreover, the allocation of ESI funds is determined based on regional eligibility criteria related to development levels, making the investment variable exogenous to current emission levels. By interacting expenditures with regional development levels, we account for heterogeneous effects while controlling for potential confounding factors. Nonetheless, one could still argue that regions with higher emissions may have received more funds, representing a case of reverse causality or omitted variable bias. To test the robustness of our identification, we implement a 2SLS estimation using structural and institutional instruments plausibly uncorrelated with current emissions.
@@ 4.1 Effect on emissions trend
Table 2 presents the estimation results of Eq. (3) using as the dependent variable the trends for GHGE per capita or GHGE intensity. The key independent variables are low-carbon expenditures per capita from the 2007-2013 and 2014-2020 programmes, each introduced with a five-year lag. The heterogeneous impacts of regional development levels are captured by interacting these investment variables with dummy variables for transition (TER) and less-developed regions (LDR), using developed regions (DER) as the reference category.
@@
| GHGE per capita | GHGE intensity | |||||
| (1t) | (2t) | (3t) | (4t) | (5t) | (6t) | |
| Low Carbon Exp. pc | 0.039 | -0.024 | 0.065 | -0.383*** | -0.073 | -0.037 |
@@
| GVAA,t-5 | 0.065* | 0.034 | ||||
| (0.034) | (0.038) | |||||
| GVAB-E,t-5 | 0.242*** | -0.042 | ||||
| (0.047) | (0.075) | |||||
| GVAF,t-5 | -0.135*** | -0.153*** | ||||
| (0.041) | (0.057) | |||||
| GVAG-J,t-5 | 0.129 | -0.200 | ||||
| (0.091) | (0.132) | |||||
| GVAK-N,t-5 | -0.362*** | -0.391*** | ||||
| (0.084) | (0.100) | |||||
| GVAO-U,t-5 | 0.170 | 0.310** | ||||
| (0.108) | (0.134) | |||||
| Observations | 3,872 | 3,408 | 3,408 | 3,872 | 3,408 | 3,408 |
| Notes: Robust standard errors clustered at the NUTS 2 level are reported in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. | ||||||
The estimates for per capita emissions (columns 1t–3t) show that in developed regions, the reference group, the coefficients on low-carbon expenditures are small and not statistically significant across all models. In contrast, the interaction terms for less developed regions are consistently positive and significant for the 2007–2013 programme, suggesting that greater expenditure is associated with increases in long-run emissions per capita. The interaction terms for the 2014–2020 period are also positive, although they are significant only in models 1t and 2t. For transition regions, the interaction coefficients are negative in most cases, but not statistically significant, except marginally in model 1t for 2007–2013 expenditures.
Regarding emissions intensity (columns 4t-6t), we observe a statistically significant negative effect of 2007–2013 expenditures in developed regions only in model 4t, suggesting improved emissions efficiency. However, for less developed regions, the corresponding interaction terms are again positive and significant in model 4t. As in the GHGEpc models, results for transition regions remain statistically not significant across all models.
Control variables, particularly GDP and sectoral GVA, exhibit expected significant relationships with emissions trends. Higher GDP is linked positively with emissions per capita but negatively with emissions intensity, highlighting the scale versus efficiency effects inherent in emissions analysis. Sectoral GVA variables reveal heterogeneity across economic activities. For instance, GVA in sectors B–E (industry) is positively correlated with per capita emissions, while GVA in services (sectors K–N) shows a significant negative association with both emissions indicators. These results underline the importance of regional economic structure in shaping emissions trajectories.
As regards institutional and political controls, we find that environmental policy stringency and political ideology show no statistically significant effects. The small positive coefficients on EPS suggest that national-level measures may not strongly influence regional emissions trends in per capita or intensity terms. Similarly, political orientation, as proxied by the left-right index, does not exhibit consistent effects.
4.2 Effect in the emissions cycle
Table 3 presents the estimation results of Eq. (3), using the cyclical components of GHG emissions per capita and emissions intensity as dependent variables.
| Dependent variable: | ||||||
| GHGE per capita | GHGE intensity | |||||
| (1c) | (2c) | (3c) | (4c) | (5c) | (6c) | |
| Low Carbon Exp. pc | -0.017 | -0.055** | -0.065** | 0.027 | 0.004 | -0.002 |
| (0.021) | (0.028) | (0.027) | (0.039) | (0.030) | (0.030) | |
| Low Carbon Exp. pc | -0.128 | -0.189 | -0.200 | -0.092 | -0.168 | -0.180 |
| (0.230) | (0.235) | (0.206) | (0.249) | (0.234) | (0.224) | |
| Low Carbon Exp. pcTER | -0.166 | -0.130 | -0.131 | -0.107 | 0.048 | 0.043 |
| (0.139) | (0.117) | (0.114) | (0.173) | (0.088) | (0.093) | |
| Low Carbon Exp. pcTER | 0.079 | -0.096 | -0.077 | -0.358 | -0.664 | -0.662 |
| (0.746) | (0.738) | (0.704) | (0.817) | (0.720) | (0.725) | |
| Low Carbon Exp. pcLDR | -0.100** | -0.110** | -0.130*** | -0.078 | -0.096** | -0.104** |
| (0.050) | (0.048) | (0.047) | (0.054) | (0.042) | (0.041) | |
| Low Carbon Exp. pcLDR | 0.198 | 0.292 | 0.333 | 0.084 | 0.221 | 0.239 |
| (0.238) | (0.240) | (0.217) | (0.253) | (0.234) | (0.224) | |
| GDPt-5 | -0.251*** | 0.082** | ||||
| (0.037) | (0.035) | |||||
| EPSt-5 | 0.010*** | 0.011*** | 0.012*** | 0.012*** | ||